OpenStudy (anonymous):

Need help It takes 1/8 tsp of baking soda to make 4 dozen cookies. How many teaspoons of baking soda does it take to make 1 dozen cookies? 1/6?

OpenStudy (anonymous):

Ok, 1/8 over 48 ?

OpenStudy (photon336):

\[\frac{ 1/8~tsp }{ 4~dozen~cookies } = \frac{ x~tsp}{ 1~dozen~cookies}\]

OpenStudy (photon336):

do you know why It can be set up this way?

OpenStudy (anonymous):

Because we want to get the unit rate?

OpenStudy (anonymous):

Yep, hold on

OpenStudy (photon336):

okay, show us your work too

OpenStudy (anonymous):

 @Photon336